OpenStudy (anonymous):

how many people you need in a room to have a probability > 0.5 of three ppl with same birthday (just day and month excluding year)

OpenStudy (anonymous):

so 366 days right

OpenStudy (anonymous):

would i just use bernoullis equation?

OpenStudy (wolf1728):

Wikipedia has a good explanation: http://en.wikipedia.org/wiki/Birthday_problem

OpenStudy (anonymous):

reviewing still no luck thou

OpenStudy (anonymous):

so 1- pn where pn= (369-n/366)

OpenStudy (wolf1728):

The answer is with 23 people, the probability is greater than 50% that 2 people will have the same birthday. Want the calculations ? Okay.

OpenStudy (anonymous):

for at least three people not just 2

OpenStudy (wolf1728):

probability of 3 people NOT having the same birthday = p(n) = 1 * (1- (1/365)) * (1- (2/365)) = 0.9917958341 (Not exactly sure if I got those numbers correct)

OpenStudy (anonymous):

so the question was how many ppl are needed in order to have at least 3 ppl with the same birthday,

OpenStudy (wolf1728):

Wikipedia's explanation is not too clear I'll have to look at it some more so now it's 3 people?

OpenStudy (anonymous):

lol the question was from the start with 3 people if not i would have figured this out long time ago

OpenStudy (wolf1728):

okay - I'll see what I can find

OpenStudy (anonymous):

so i researched it this looks pretty complicated, now trying to figure out how they derived such formula

OpenStudy (anonymous):

see if you can make some sense on it

OpenStudy (anonymous):

ill just use there method for now lol

OpenStudy (wolf1728):

I just got a phone call - I'll be a while

OpenStudy (wolf1728):

I found this at the Online Encyclopedia of Integer Sequences: http://oeis.org/A014088 The minimal number of people you will need in order to have a better than 50% chance of 3 people having the same birthday is 88 and if you go to this page: http://oeis.org/A014088/a014088.txt you will see that the probability = 0.499454850632 for 3 people having the same birthday out of a group of 87 people and probability = 0.511065110625 when increasing the group to 88 people. I couldn't find a formula for this, so I hope what you found at stackexchange.com will help you compute this. Anyway, I am definitely sure that the number of people you would need is 88.